Ph of a mixture of 1 M benzoic acid (pKapKa = 4.20) and 1 M solution benzoate is 4.5. In 300 mL of buffer solution, calculate the moles of benzoic acid in 300 ml bufffer

Grade:11

1 Answers

Siddharth Kushwaha
5 years ago
pH Formula of acidic buffer
 
pH = pKa + log(C/A)
 
4.5=4.2 + log(1/A)
 
Antilog(0.3) = 1/A
 
A = 0.50 M
 
MV = 0.50×0.3
 
= 0.15 moles